OpenStudy (anonymous):

What is the solution of log(2x + 6) 8 = 3

OpenStudy (anonymous):

x = −2 x = 2 x = −3 x = 3

OpenStudy (bibby):

(2x+6)^3 = 8 hint: you can write 8 as 2^3

OpenStudy (ranga):

\[\Large \log_{b} A = C \quad means \quad A = b^C\]

OpenStudy (anonymous):

is it -2?

OpenStudy (ranga):

Yes, x = -2

OpenStudy (anonymous):

yay i tries solving it the way you guys all taught me :)

OpenStudy (anonymous):

i need help with others

OpenStudy (ranga):

Post the questions and hopefully someone will be able to answer.

OpenStudy (anonymous):

For the graphed function f(x) = (4)^(x - 1) + 2, calculate the average rate of change from x = 2 to x = 4.

OpenStudy (anonymous):

is there a formula for this?

OpenStudy (ranga):

Average rate of change of f(x) over the interval [a, b] = { f(b) - f(a) } / (b - a)

OpenStudy (anonymous):

soo.. how would i do that

OpenStudy (ranga):

ou are given f(x). Here the interval is [2, 4]. So a = 2, b = 4 Find f(4) by setting x = 4 in the function. Find f(2) by setting x = 2 in the function. Plug the numbers into the formula for average rate of change.

OpenStudy (anonymous):

f(4)-f(a)/4-2?

OpenStudy (ranga):

{ f(4) - f(2) } / (4 - 2)

OpenStudy (anonymous):

2**

OpenStudy (anonymous):

so that would give me..

OpenStudy (ranga):

What are you getting?

OpenStudy (anonymous):

-30?

OpenStudy (anonymous):

68

OpenStudy (ranga):

Try again.

OpenStudy (ranga):

Is this the function? \[\Large f(x) = 4^{x-1} + 2\]

OpenStudy (anonymous):

yes

OpenStudy (ranga):

f(4) = 4^(4-1) + 2 = 4^3 + 2 = 64 + 2 = 66

OpenStudy (anonymous):

thats the solution? 66 wow thank you

OpenStudy (anonymous):

thats it?

OpenStudy (ranga):

No 66 is not the solution. I just evaluated it at x = 4. You can evaluate it at x = 2 in a similar way. Then find { f(4) - f(2) } / (4 - 2)

OpenStudy (anonymous):

i see, i got 8.

OpenStudy (anonymous):

and then 58/2...

OpenStudy (ranga):

f(4) = 4^(4-1) + 2 = 4^3 + 2 = 64 + 2 = 66 f(2) = 4^(2-1) + 2 = 4^1 + 2 = 6 Average rate of change of f(x) over the interval [2, 4] = { f(4) - f(2) } / (4 - 2) = (66-6)/(4-2) = 60/2 = 30

OpenStudy (anonymous):

ohh, wow i had 8... careless mistakes. this was really helpful :) thank you

OpenStudy (ranga):

you are welcome.